While shopping, you find that a 12-oz package of brand A costs $2.40 and a 16-oz package of brand B costs $2.50. Which brand has a lower unit cost?
Answer by srinivas.g(540)

12 -oz package of brand A cost = $2.40
1 unit cost = 2.40 /12
= $ 0.2
a 16-oz package of brand B costs $2.50
1 unit cost = 2.50/16
=$ 0.15
Brand B has lower unit cost